# Gauss Principle in Incompressible Flow: A Unified Variational Perspective on Pressure and Projection

### v1

@@ 3.3 Uniqueness of Dirichlet Orthogonality
Peters & Ormiston (2025) address the above arbitrariness by proposing an intuitive orthogonal decomposition
$\int_{\Omega}\nabla P_{R}\cdot\nabla P_{F}\,dV=0$ (6)
@@
3.4 Physical choice of impressed subspace
For the exterior airfoil problem the physical implications are immediate. If $S_{\rm imp}$ contains the potential-flow circulation mode, taking $P_{F}=\Gamma\,\psi_{s}$ neatly encodes the circulation input while leaving $P_{R}$ to enforce the constraints for the actual frozen velocity. However, we now show that even with Dirichlet orthogonality, the choice of impressed subspace is highly arbitrary.
@@ 3.4 Physical choice of impressed subspace
    Circulation basis . Choose $\mathcal{S}_{\!\rm imp}=\mathrm{span}\{\psi_{s}\}$ so that $\nabla\psi_{s}$ represents the potential–flow circulation mode (the harmonic field tangent to the wall with unit loop integral). Then $P_{F}=\Gamma\,\psi_{s}$ carries precisely the circulation degree of freedom emphasized in Gonzalez & Taha (2022), and $P_{R}$ enforces the constraints. This matches the “circulation‑aware” split described in  Peters & Ormiston (2025).
@@
    Freestream split. It is also legitimate to designate the freestream direction as impressed. A minimal way is to include the linear potential $\phi_{\infty}(x)=U_{\infty}\!\cdot x$ among the generators and let $\psi_{\infty}:=\rho\,\phi_{\infty}$. Then $\nabla P_{F}$ may contain the constant vector $\rho U_{\infty}$. This merely tilts the “target” in $C=(u\!\cdot\!\nabla)u+\nabla P_{F}/\rho$; the Poisson step automatically compensates so that $u_{t}$ remains divergence‑free and wall‑tangent. There is no requirement that this impressed piece satisfy the wall condition—it need not, because only $u_{t}$ is constrained.
@@
    Full potential reference. One may also place the entire steady Bernoulli gradient for a chosen reference potential flow into $P_{F}$ (e.g. the zero‑circulation solution $u_{0}$ or $u_{0}+\Gamma u_{1}$). In steady irrotational settings, $(u\!\cdot\!\nabla)u=\nabla(\tfrac{1}{2}|u|^{2})$; taking $P_{F}=\rho\,\tfrac{1}{2}|u_{\rm ref}|^{2}$ pushes that Bernoulli content into the impressed side and leaves $P_{R}$ to enforce the constraints for the actual frozen $u$.
The latter two results are particularly important given the criticism of  Peters & Ormiston (2025) of  Gonzalez & Taha (2022)’s derivation as incorrect because they ignored impressed pressure. While in some cases (e.g. accounting for body forces due to gravity), impressed forces can be clearly delineated, it is indeed not incorrect to consider the entire field as non-impressed in the typical airfoil scenario. In other words, even if one impresses the uniform stream, or even a full steady potential reference, the same mechanism applies: the Gauss step removes only what is needed to satisfy the constraints and leaves the impressed content intact. In all cases the reaction is irrotational and hence cannot create or destroy vorticity; it instantaneously redistributes pressure to maintain solenoidality and wall compatibility while vorticity evolves through advection (and any impressed physics).
Finally, this viewpoint unifies interpretation and practice. The standard predict–Poisson–correct algorithm is precisely the Gauss–Appell step on a grid, and $\mathcal{A}^{\star}$ furnishes a diagnostic that can be monitored in computation: spikes signal incompatible boundary data, under-resolved features, or an unhelpful choice of impressed content. In reduced-order or data-assisted settings, the choice of $S_{\rm imp}$ allows trusted physics to be injected while the same projection guarantees exact enforcement of incompressibility and wall conditions.
@@ Acknowledgments
Substitute the correction 3.3 into the orthogonality condition 6:
@@
Peters & Ormiston (2025) cast multipliers via a finite‑dimensional Lagrange‑multiplier theorem, tying the number of independent multipliers to the rank of a constraint Jacobian. That language is useful but does not transfer cleanly to the PDE setting of incompressible flow, where the constraint is an operator ($\nabla\!\cdot$) and the multiplier $P_{R}$ is an entire field. What replaces “rank” is the orthogonal projection onto the divergence‑free, wall‑tangent subspace (Leray projector), with the multiplier recovered as the gradient part in the Helmholtz–Hodge decomposition; see the Euler/Appellian equivalence proof in Gonzalez & Taha (2022). In other words, pressure acts as the constraint that removes the non‑solenoidal content of $\rho(u_{t}+C)$, and the correct mathematical object is a projection, and not a finite matrix rank.
Appendix C: Lagrangian form
In contrast to prior literature, we additionally present the derivation in Lagrangian form, as it is more intuitive and applies more naturally to the steady state treatment. Let $\Omega_{0}$ be a reference domain with boundary $\partial\Omega_{0}$. Consider instantaenous particle locations $X(\xi,\eta,t)$, where $\xi,\eta$ are reference coordinates. At the fixed time $t_{0}$, abbreviate
$X(\xi,\eta)\triangleq X(\xi,\eta,t_{0}),\qquad\dot{X}(\xi,\eta)\triangleq\partial_{t}X(\xi,\eta,t_{0}),\qquad a(\xi,\eta)\triangleq\ddot{X}(\xi,\eta,t_{0})\in\mathbb{R}^{2}.$
Kinematic relations:
$J=X_{\xi}Y_{\eta}-X_{\eta}Y_{\xi},\qquad\widetilde{J}=\begin{pmatrix}Y_{\eta}&-\,Y_{\xi}\\ -\,X_{\eta}&X_{\xi}\end{pmatrix},\qquad\partial_{\xi}\widetilde{J}_{i\xi}+\partial_{\eta}\widetilde{J}_{i\eta}=0\ (i=1,2).$ (11)
Time derivatives:
$\dot{J}=\widetilde{J}_{ij}\,\partial_{x_{0}j}\dot{X}_{i},\qquad\ddot{J}=\widetilde{J}_{ij}\,\partial_{x_{0}j}a_{i}+\dot{\widetilde{J}}_{ij}\,\partial_{x_{0}j}\dot{X}_{i},\qquad\dot{\widetilde{J}}=\begin{pmatrix}\dot{Y}_{\eta}&-\,\dot{Y}_{\xi}\\ -\,\dot{X}_{\eta}&\dot{X}_{\xi}\end{pmatrix}$. (12)
Feasibility assumptions at $t_{0}$ (hold before the Gauss step):
  • Incompressibility: $J(\xi,\eta)=1$ in $\Omega_{0}$.
  • Velocity compatibility: $\dot{J}(\xi,\eta)=\widetilde{J}_{ij}\,\partial_{x_{0}j}\dot{X}_{i}=0$ in $\Omega_{0}$.
  • Wall non–penetration: $\dot{X}\!\cdot(\widetilde{J}n_{0})=0$ on $\Omega_{0}$.
As mentioned previously, Gauss/Appell is instantaneous. The decision variable is $a=\ddot{X}$; $X$ and $\dot{X}$ are frozen data. Thus feasibility (position/velocity level) is assumed, not solved for, at $t_{0}$. Differentiating the feasibility relations in time and evaluate at $t_{0}$:
$\widetilde{J}_{ij}\,\partial_{x_{0}j}a_{i}\;=\;-\,\dot{\widetilde{J}}_{ij}\,\partial_{x_{0}j}\dot{X}_{i}\qquad\text{in }\Omega_{0},$ (13)
$a\!\cdot(\widetilde{J}n_{0})\;+\;\dot{X}\!\cdot(\dot{\widetilde{J}}\,n_{0})\;=\;0\qquad\text{on }\partial\Omega_{0},$ (14)
which reduces to $a\!\cdot(\widetilde{J}n_{0})=0$ in the far-field when $\dot{\widetilde{J}}\to 0$.
Least–constraint
Given an impressed material acceleration $b(\xi,\eta)$, define
$\mathcal{G}[a]\;=\;\frac{1}{2}\int_{\Omega_{0}}\rho\,|a-b|^{2}\,dA,\qquad\text{with }X,\ \dot{X}\ \text{frozen at }t_{0}.$ (15)
We introduce the interior multiplier $\lambda(\xi,\eta)$ for (13) and boundary multipliers $\mu$ on $\partial\Omega_{0}$. (We do not add multipliers for $J=1$, $\dot{J}=0$, $u\!\cdot n=0$, since those are feasibility of $(X,\dot{X})$.)
$\mathcal{L}[a,\lambda,\mu_{w},\mu_{\infty}]=\int_{\Omega_{0}}\Big\{\tfrac{1}{2}\,\rho\,|a-b|^{2}+\lambda\big(\widetilde{J}_{ij}\partial_{x_{0}j}a_{i}+\dot{\widetilde{J}}_{ij}\partial_{x_{0}j}\dot{X}_{i}\big)\Big\}\,dA+\int_{\partial\Omega_{0}}\mu\big(a\!\cdot(\widetilde{J}n_{0})+\dot{X}\!\cdot(\dot{\widetilde{J}}n_{0})\big)\,ds_{0}.$ (16)
Write the interior term explicitly:
$\widetilde{J}_{ij}\,\partial_{x_{0}j}a_{i}=Y_{\eta}\,a_{x,\xi}-Y_{\xi}\,a_{x,\eta}-X_{\eta}\,a_{y,\xi}+X_{\xi}\,a_{y,\eta}.$
Applying the first variation and spatial integration by parts in $\xi$ and $\eta$ gives the interior Euler–Lagrange equations
${\begin{aligned} \rho(a_{x}-b_{x})-(\lambda Y_{\eta})_{\xi}+(\lambda Y_{\xi})_{\eta}&=0,\\ \rho(a_{y}-b_{y})+(\lambda X_{\eta})_{\xi}-(\lambda X_{\xi})_{\eta}&=0,\end{aligned}}\qquad\text{in }\Omega_{0},$ (17)
and boundary compatibility $\quad\lambda+\mu=0\quad\text{on }\partial\Omega_{0}\quad$}.
$\delta\lambda\ \ :\ \ {\quad Y_{\eta}a_{x,\xi}-Y_{\xi}a_{x,\eta}-X_{\eta}a_{y,\xi}+X_{\xi}a_{y,\eta}=-\,\dot{\widetilde{J}}_{ij}\partial_{x_{0}j}\dot{X}_{i}\quad\text{in }\Omega_{0}.\quad}$ (18)
$\delta\mu:\ \ {\quad a\!\cdot(\widetilde{J}n_{0})+\dot{X}\!\cdot(\dot{\widetilde{J}}n_{0})=0\quad\text{on }\partial\Omega_{0}\quad}$ (19)
Thus we have, for feasibility at $t_{0}:$ $J=1,\qquad\widetilde{J}_{ij}\partial_{x_{0}j}\dot{X}_{i}=0\quad\text{in }\Omega_{0},\qquad\dot{X}\!\cdot(\widetilde{J}n_{0})=0\quad\text{on }\partial\Omega_{0}.$ Therefore,
(acceleration level by Gauss)
$\widetilde{J}_{ij}\partial_{x_{0}j}a_{i}$ $=-\,\dot{\widetilde{J}}_{ij}\partial_{x_{0}j}\dot{X}_{i}$ $\text{in }\Omega_{0},$ (20)
$a\!\cdot(\widetilde{J}n_{0})+\dot{X}\!\cdot(\dot{\widetilde{J}}n_{0})$ $=0$ $\text{on }\partial\Omega_{0}.$ (21)
[…]
$\lambda+\mu$ $=0$ $\text{on }\partial\Omega_{0}.$ (23)
Transforming to Eulerian Coordinates: It is more intuitive to write these expressions in the Eulerian form: Let $x=X(\xi,\eta,t_{0})$, $J=1$, $u=\dot{X}^{\sharp}$, $a=\ddot{X}^{\sharp}=D_{t}u$, $P_{R}=\lambda^{\sharp}$, and $b^{\sharp}=b\circ X^{-1}$. We will use Piola/Nanson transformations (Gurtin, 2010):
$\operatorname{Div}_{0}(\lambda\,\widetilde{J})\mapsto\nabla_{x}P_{R},\qquad\widetilde{J}_{ij}\partial_{x_{0}j}(\cdot)\mapsto\nabla_{x}\!\cdot(\cdot)^{\sharp},\qquad(\widetilde{J}n_{0})\,ds_{0}\mapsto n\,ds,$
where the push forward transformation $\mapsto$ and superscript $(\cdot)^{\sharp}$ denote the transfer from reference coordinates to current coordinates.
(feasible at $t_{0}$)
$\nabla_{x}\!\cdot u=0\ \text{ in }\Omega,\qquad u\!\cdot n=0\ \text{ on }\partial\Omega_{0}.$ (24)
(acceleration level)
$\nabla_{x}\!\cdot(D_{t}u)$ $=\nabla_{x}\!\cdot\big((u\!\cdot\!\nabla_{x})u\big)$ $\text{in }\Omega,$ (25)
$D_{t}u\!\cdot n$ $=(u\!\cdot\!\nabla_{x}u)\!\cdot n$ $\text{on }\partial\Omega_{0}.$ (26)
$\rho\,(D_{t}u-b^{\sharp})+\nabla_{x}P_{R}$ $=0$ $\text{in }\Omega,.$ (27)
If $b^{\sharp}=-(1/\rho)\nabla_{x}P_{F}$, then $p=P_{F}+P_{R}$ and $\rho(\partial_{t}u+(u\!\cdot\!\nabla)u)=-\nabla_{x}p$.
Given (Lagrangian wall constraint at the instant $t_{0}$)
$a\cdot(\widetilde{J}\,n_{0})\;+\;\dot{X}\cdot\big(\dot{\widetilde{J}}\,n_{0}\big)\;=\;0.$ (28)
We use two identities Gurtin (2010) in the push–forward. At $t_{0}$:
$\text{(Nanson)}\qquad\widetilde{J}\,n_{0}\,dS_{0}\;=\;J\,n\,dS,$ (29)
$\text{(cofactor time derivative)}\qquad\dot{\widetilde{J}}\;=\;-\,(\nabla_{x}u)^{\!T}\,\widetilde{J}$. (30)
Multiply (28) by the reference line element $dS_{0}$ and use (29)–(30):
$\big[a\cdot(\widetilde{J}n_{0})\big]\,dS_{0}\;=\;a\cdot n\,dS,$
$\big[\dot{X}\cdot(\dot{\widetilde{J}}n_{0})\big]\,dS_{0}\;=\;u\cdot\big(-(\nabla_{x}u)^{\!T}\widetilde{J}n_{0}\big)\,dS_{0}\;=\;-\,u\cdot\big((\nabla_{x}u)^{\!T}n\big)\,dS,$
$\textrm{Hence }\big(a\cdot n-u\cdot[(\nabla_{x}u)^{\!T}n]\big)\,dS=0,\qquad\Rightarrow\qquad a\cdot n\;=\;u\cdot\big[(\nabla_{x}u)^{\!T}n\big]\ \ \text{on the wall.}\,$
Because $a=\partial_{t}u+(u\!\cdot\!\nabla_{x})u$ and $(\nabla_{x}u)\,u=(u\!\cdot\!\nabla_{x})u$,
$u\cdot[(\nabla_{x}u)^{\!T}n]\;=\;\big((\nabla_{x}u)\,u\big)\cdot n\;=\;(u\!\cdot\!\nabla_{x}u)\cdot n,$
so $(\ast)$ is equivalent to
$(\partial_{t}u+(u\!\cdot\!\nabla_{x})u)\cdot n\;=\;(u\!\cdot\!\nabla_{x}u)\cdot n\rightarrow u_{t}\cdot n=0.$

### v2

@@ 3.4 Uniqueness of Dirichlet Orthogonality
Peters & Ormiston (2025) propose the following Dirichlet orthogonality as a convenient gauge-fixing convention for selecting a unique representative of the bookkeeping split $p=P_{F}+P_{R}$ (once a reference/impressed subspace has been chosen):
$\int_{\Omega}\nabla P_{R}\cdot\nabla P_{F}\,dV=0.$ (6)
It is stressed that (6) is not a physical restriction on an impressed force defined by external physics; it is a convention that fixes the gauge in the pressure bookkeeping. If $P_{F}$ is reserved for physically prescribed impressed potentials, the split is already unique (with $P_{R}$ the unique Lagrange multiplier enforcing the constraints) and (6) need not hold.
@@
3.5 Physical impressed forces vs. reference (gauge) choices
For the exterior airfoil problem the distinction between physical impressed forces and bookkeeping/reference choices is immediate. The circulation degree of freedom in exterior potential flow resides in the harmonic subspace of the velocity (divergence-free, curl-free, wall-tangent in a multiply connected domain) and cannot be represented as $\nabla\psi$ of a single-valued scalar pressure potential; hence it cannot be embedded into $P_{F}$ as a physical pressure field. By contrast, conservative body-force potentials (e.g. gravity) may legitimately be placed in $P_{F}$ as prescribed impressed physics. If one introduces additional reference content in $P_{F}$ for diagnostics, then that choice is conventional.
@@
    Harmonic velocity modes (circulation) are not pressure gradients. In an exterior domain, circulation enters through the harmonic subspace of the velocity (divergence-free, curl-free, wall-tangent with nonzero loop integral). Such modes are not of the form $\nabla\psi$ for any single-valued scalar field, so they cannot be represented as a pressure potential in $P_{F}$ (nor in the reaction pressure $P_{R}$). In the fixed-time Gauss/Appell step, the circulation parameter $\Gamma$ appears only through the chosen frozen state $u(\cdot,t)$.
@@
    Conservative body-force potential. It is legitimate to include a prescribed uniform acceleration (e.g. gravity) through an impressed potential $P_{F}(x)=-\rho\,g\cdot x$, for which $(1/\rho)\nabla P_{F}=-g$ has units of acceleration. The Poisson step then returns $P_{R}$ to enforce $\nabla\cdot u_{t}=0$ and $u_{t}\cdot n=0$ for the frozen $u$.
@@
    Reference pressure for diagnostics (gauge). One may define a reference pressure for interpretation, e.g. subtract the Bernoulli pressure of a chosen reference potential flow in post-processing: $p^{\prime}:=p+\tfrac{1}{2}\rho|u_{\rm ref}|^{2}$. This is a change of variables (gauge) and should not be interpreted as an impressed force. For a given snapshot $u$ with no impressed forces, the pressure gradient is entirely the reaction that enforces incompressibility and wall compatibility (the Leray projection of $\rho(u\cdot\nabla)u$).
These remarks are particularly relevant to the discussion in Peters & Ormiston (2025) regarding “impressed pressure”. In the present framework, the frozen-time Gauss/Appell step returns
$u_{t}^{\star}=-\Big((u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla P_{F}+\frac{1}{\rho}\nabla P_{R}\Big),$
where $P_{F}$ is reserved for prescribed conservative impressed forcing (if any), and $P_{R}$ is the reaction (constraint) pressure. Enforcing the acceleration-level constraints $\nabla\!\cdot u_{t}^{\star}=0$ in $\Omega$ and $u_{t}^{\star}\!\cdot n=0$ on $\partial\Omega$ yields the Poisson–Neumann problem
$\nabla^{2}P_{R}=-\rho\,\nabla\!\cdot\Big((u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla P_{F}\Big),\qquad\partial_{n}P_{R}=-\rho\Big((u\!\cdot\!\nabla)u+\frac{1}{\rho}\nabla P_{F}\Big)\!\cdot n,$
which determines $P_{R}$ uniquely up to an additive constant. A genuinely impressed pressure potential corresponds to an externally prescribed conservative force $f=-\nabla\Phi$, for which one may take $P_{F}=\rho\Phi$; in the unforced exterior Euler airfoil setting $P_{F}\equiv 0$ and the entire pressure is reaction/constraint pressure.
Additionally, since $\nabla\times\nabla(P_{R}+P_{F})=0$, the pressure reaction is irrotational and thus does not appear explicitly in the interior vorticity evolution obtained by taking the curl of the momentum equation: $\omega_{t}+(u\!\cdot\!\nabla)\omega=(\omega\!\cdot\!\nabla)u$ (for conservative impressed forcing). We note that this statement concerns the bulk equation; boundary vorticity fluxes can be expressed in terms of tangential pressure gradients (e.g. Lighthill (1963) and Morton (1984).)
Finally, this viewpoint unifies interpretation and practice. The standard predict–Poisson–correct algorithm is precisely the Gauss–Appell step on a grid, and $\mathcal{A}^{\star}$ furnishes a diagnostic that can be monitored in computation: spikes may signal incompatible boundary data, under-resolved features, or an unhelpful choice of impressed/reference content. In reduced-order or data-assisted settings, the choice of $S_{\rm imp}$ allows trusted physics to be injected while the same projection guarantees exact enforcement of incompressibility and wall conditions.
@@ Acknowledgments
Substitute the correction 3.4 into the orthogonality condition 6:
@@
Peters & Ormiston (2025) cast multipliers via a finite-dimensional Lagrange-multiplier theorem, tying the number of independent multipliers to the rank of a constraint Jacobian. In a discrete setting this perspective can be helpful. In the PDE setting, however, incompressibility and impermeability impose pointwise constraints. The Leray projector extracts the divergence-free, wall-tangent component, and the reaction pressure is recovered from the gradient component in the decomposition. The velocity space also contains a finite-dimensional harmonic subspace (e.g. circulation modes), which is left unchanged by the projector; see e.g. Sohr (2012) and the Euler/Appellian equivalence proof in Gonzalez & Taha (2022).
This viewpoint also makes the analogue of a “rank/consistency” condition explicit. The Neumann Poisson problem for $P_{R}$ is solvable if and only if the volume source and boundary flux satisfy the compatibility condition,
$\int_{\Omega}(-\,\nabla\!\cdot C)\,dV=\int_{\partial\Omega}(-\,C\!\cdot n)\,dS,$
which holds identically by the divergence theorem. Uniqueness is up to an arbitrary constant.
Physically, the decomposition expresses the familiar fact that pressure is a reaction field that enforces ideal kinematic constraints without doing work on admissible virtual velocities. Indeed, for any $v\in V$,
$\int_{\Omega}v\cdot\nabla P_{R}\,dV=\int_{\partial\Omega}P_{R}\,v\cdot n\,dS-\int_{\Omega}P_{R}\,\nabla\!\cdot v\,dV=0.$
Thus the reaction pressure is $L^{2}$-orthogonal to the space of admissible velocities, exactly as in d’Alembert’s principle, and the “multiplier” is most naturally interpreted as the gradient component required to project $\rho(u_{t}+C)$, rather than as a finite-dimensional rank count.